OpenStudy (anonymous):

Given the equation y − 4 = three fourths(x + 8) in point-slope form, identify the equation of the same line in slope-intercept form. −three fourthsx + y = 10 3x − 4y = −40 y = three fourthsx + 12 y = three fourthsx + 10

OpenStudy (anonymous):

\[y − 4 = \frac{3}{4}(x + 8) \rightarrow 4(y − 4) = 3(x + 8)\] \[ \rightarrow 4y − 16 = 3x + 24 \rightarrow 4y = 3x + 24+16\] \[ \rightarrow 4y = 3x + 40 \rightarrow y = \frac{3}{4}x + \frac{40}{4}\] \[\rightarrow y = \frac{3}{4}x + 10\] Hence option " y = three fourthsx + 10 " is the correct answer. @Tayloresc

OpenStudy (texaschic101):

do you understand everything dpasangh did ?

OpenStudy (anonymous):

yes